Exercise 14-12

On January 2, 2012, Waterway Corporation issued $1,100,000 of 10% bonds at 98 due December 31, 2021. Interest on the bonds is payable annually each December 31. The discount on the bonds is also being amortized on a straight-line basis over the 10 years. (Straight-line is not materially different in effect from the preferable “interest method”.)

The bonds are callable at 101 (i.e., at 101% of face amount), and on January 2, 2017, Waterway called $660,000 face amount of the bonds and redeemed them.

Ignoring income taxes, compute the amount of loss, if any, to be recognized by Waterway as a result of retiring the $660,000 of bonds in 2017. (Round answer to 0 decimal places, e.g. 38,548.)

Loss on redemption $

  


Prepare the journal entry to record the redemption. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

January 2, 2017
Answer #1
Discount related to Bonds redeemed 13200 =660000*(1-0.98)
Unamortized Discount related to Bonds redeemed 6600 =13200*5/10
1
Bonds redemption price 666600 =660000*1.01
Less: Carrying value of bonds redeemed 653400 =660000-6600
Loss on redemption 13200
2
Debit Credit
Bonds payable 660000
Loss on redemption of Bonds 13200
       Discount on Bonds payable 6600
       Cash 666600 =660000*1.01